You are an analyst employed by a yacht manufacturer that last year sold 30,000 luxury yachts at $500,000 each. Your market research indicates that:


i) the price elasticity of demand for your aircrafts in −0.5. (or +0.5 in absolute value); 

ii) the income elasticity of demand for your aircrafts is +2.6; and 

iii) the cross price elasticity for your aircrafts with respect to the price of a comparable jet manufactured by a competitor is +1.4

 Now suppose that the only change you expect is a 6% increase in the price of your aircraft. Estimate sales this year and discuss the impact on total revenues. (5 MARKS)

]



1
Expert's answer
2022-04-05T10:32:40-0400

a) The quantity sold will decrease by 6×0.5=3%, or by 30,000×0.03=900 yachts till 30,000×(1-0.03)=29,100 yachts. The revenue will increase because the price increases to a greater extent than the quantity decreases. New revenue is 29,100×500,000×1.06=15,423 million dollars, the revenue before price change is 30,000×500,000=15,000 million dollars